OpenStudy (anonymous):

Carol and Mary are the two tallest basketball players on their high school team. Carol is 69 1/3 inches tall. And Mary is 70 1/4 inches tall. How much taller is Mary than Carol?

OpenStudy (anonymous):

It is simple subtraction of fractions..... we first convert them to improper fractions as follows: Carol's height : 69 1/3 = 208/3 Mary's height : 70 1/4 = 281/4 Now we find their denominators are not same, so we find LCD of 3 and 4, which is 12 So we make both fraction to have same base as follows 208 4 832 ----* --- = ------ 3 4 12 and 281 3 843 ----*--- = ------ 4 3 12 We find 843> 832 so 843/12 > 832/12 Now subtracting 843 832 843 - 832 11 ---- - ----- = ----------- = ---- 12 12 12 12 So Mary is taller than Carol by 11/12 inch